OpenStudy (anonymous):

one over x to the 9th power equals what

OpenStudy (schrodingers_cat):

1/x^9 = x^-9

OpenStudy (anonymous):

from the rule of exponents:\[a^n=\frac{1}{a^{-n}}\]